ANTI-HYDROGEN BONDING INTERACTION IN TETRAFLUOROMETHANE-WATER

\underline{A. MARIS, W. CAMINATI}, A. DELL’ERBA, P. G. FAVERO, Dipartimento di Chimica ”G.Ciamician”, Universita’ di Bologna, Via Selmi I-40126, Italy.

The rotational spectrum of the tetrafluoromethane-water (CF_{4}-W) complex has been assigned in a supersonic expansion through millimeter wave absorption spectroscopy. In order to reach the best experimental conditions mixtures with rather high concentrations of CF_{4} (up to 20-30 per cent) in Ar at a pressure of 50 kPa were flown at room temperature and expanded to about 1 Pa through a pulsed nozzle with a 0.35 mm diameter. All measured isotopomers (CF_{4}—H_{2}O, CF_{4}—H_{2}^{18}O, CF_{4}—D_{2}O, CF_{4}—DOH) display symmetric top (prolate) spectra, because of the nearly free rotation of the water subunit with respect to CF_{4}. The first two isotopomers are split into two component lines due to an internal rotation of water along its C_{2}v axis. The water is linked to CF_{4} by a C-F—O, anti hydrogen bond, interaction (see below).